Circular No. 6650

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 H2 (SOHO) C. St. Cyr, NASA Goddard Space Flight Center, reports that Shane

Stezelberger, SOHO-LASCO Operations, has discovered a comet on images obtained with coronagraphs aboard the SOHO mission, which is located near the Lagrangean point 0.01 AU sunward of the earth. Detections were made with the C3 “outer” coronagraph during Apr. 29.3–30.0 UT, the comet moving in from 31.43 R_{⊙} from the sun’s center in p.a. 296^{o}.2 (with respect to solar north); with the C2 “inner” coronagraph during May 1.6– 1.8 and 2.5–2.6 (minimum 2.88 R_{⊙}); and with C3 during May 3.0–5.0 out to 30.51 R_{⊙} in p.a. 114^{o}.6. J2000.0 positions, reduced by G. V. Williams and given in more detail on MPEC 1997-J05, are summarized below, together with parabolic orbital elements by the undersigned. The ephemeris is uncertain, because of difficulty correlating the C2 and C3 positions.

1997 UT  α_{2000}  δ_{2000}  m_{1}
Apr. 29.265  1^{h}50.^{m}6  +14^{o}43′
30.024  2 01.8  +14 57
May 1.649  2 27.8  +15 21  2–5
1.826  2 30.6  +15 24
2.479  2 41.0  +15 35
2.597  2 42.9  +15 36
3.020  2 50.3  +15 41
4.990  3 21.7  +16 00
T = 1997 May 2.83 TT  ω = 191.^{o}57
Ω = 227.21 }2000.0
q = 0.1373 AU  i = 18.42
1997TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 10 4^{h}26.^{m}37 +15^{o}50.^{′}1 0.957 0.326 18.^{o}9 89.^{o}7  8.0
11 4 38.30 +15 45.4 0.936 0.359 20.8 91.3  8.4
12 4 50.16 +15 39.3 0.917 0.390 22.7 92.1  8.7
13 5 02.00 +15 31.6 0.901 0.421 24.6 92.4  9.0
14 5 13.84 +15 22.3 0.886 0.452 26.5 92.3  9.3
15 5 25.67 +15 11.4 0.873 0.482 28.4 91.9  9.5
16 5 37.50 +14 58.7 0.862 0.511 30.4 91.1  9.8
17 5 49.30 +14 44.3 0.854 0.540 32.3 90.2 10.0
18 6 01.06 +14 28.1 0.847 0.568 34.2 89.0 10.2
19 6 12.75 +14 10.2 0.842 0.596 36.1 87.7 10.4
20 6 24.34 +13 50.7 0.838 0.624 37.9 86.3 10.6
1997 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden